# #EpiTwitter: Public Health Messaging During the COVID-19 Pandemic

### v1

@@ Results
The results presented in SI Table LABEL:tab:perspec demonstrate the contrasting viewpoints between the two groups on various issues. Regarding the origins of the virus, PHEs generally lean towards the belief that it originated in a laboratory, albeit with some skepticism, while pseudo-experts heavily criticize China and its potential involvement in gain of function research. PHEs emphasize the importance of ongoing vigilance, adherence to stay-at-home orders, and widespread use of masks, whereas pseudo-experts question the effectiveness of lockdowns and mask mandates and criticize government intervention in these areas. On the topic of therapeutics, PHEs urge caution against self-prescribing drugs like Hydroxychloroquine, Azithromycin, and Ivermectin without evidence of their efficacy in treating COVID-19, whereas pseudo-experts advocate for the use of these medications.
@@
Emotions fluctuate over time and in response to events. Figure 6 illustrates the temporal dynamics of positive emotions expressed by PHEs and pseudo-experts. We leverage ChatGPT to summarize changes in emotions expressed and are discussed further in SI Table LABEL:tab:emotions_summ. Optimism and joy among PHEs surge following the announcement of stay-at-home orders post March 15, 2020. This can be attributed to factors such as gratitude for guidance by then-New York Governor Andrew Cuomo and enhanced accessibility to COVID-19 testing. Similarly, we note a corresponding albeit smaller increase among pseudo-experts, particularly in response to President Trump’s management of the pandemic and France’s endorsement of HCQ as a viable COVID-19 treatment.
@@
We also examine the use of moral language by the two groups. SI Figure 13 compares the distribution of the daily share of tweets expressing each moral foundation. Overall, PHEs use more positive moral language, emphasizing the dimensions of care, fairness, authority, loyalty and purity, while pseudo-experts tend to prefer the negative moral dimensions of harm, cheating, subversion and betrayal. The differences in use of moral language are more subdued compared to those for emotions. SI Figure 14 illustrates the temporal dynamics of positive and negative moral language used by PHEs and pseudo-experts. We summarize the positive spikes using ChatGPT in SI Table LABEL:tab:mft_summ.
@@
Figure 7 compares the log-odds along with the corresponding standard errors of estimation, to show which group used more emotional language to frame a specific issue. The plot highlights differences between the two groups and gives insights into emotionally charged issues. The biggest gap in emotions appears on the issue of lockdowns, where pseudo-experts are far more likely to express anger, disgust and sadness than PHEs. This position is consistent with the efforts to end the lockdowns (see Great Barrington declaration Burki (2021)). The second largest gap in emotions appears in the discussion of therapeutics, where PHEs are more likely to express anger and disgust, but pseudo-experts are less likely to use these emotions. Pseudo-experts also use more positive language with more joy and optimism when talking about therapeutics, in contrast to PHEs, consistent with the highly contentious debates about this issue. Other notable differences include pseudo-experts expressing more fear and less joy about vaccines, while PHEs express less fear and more optimism.
@@
Studies show that public response to the pandemic became polarized fairly quickly, with political partisanship shaping online activity and discussions about the pandemic already in the early stages of the pandemic Jiang et al. (2020); Rao et al. (2021). Online discussions also became emotionally polarized: when interacting with members of the opposite party, Twitter users expressed more anger and disgust, more toxicity, and less joy than in their interactions with same-party members Feldman et al. (2023). Such interactions are characteristic of affective polarization Iyengar and Westwood (2015), patterns of in-group love and out-group hate that have contributed to the growing partisan divide and the erosion of trust between the two parties in US. As a result, partisanship predicted the adoption of pandemic prevention measures more than other factors Grossman et al. (2020).
@@
Figure 15 lists the top ten accounts that are more likely to be mentioned by PHEs and pseudo-experts positively or negatively. We then assess the retweet interactions of public health and pseudo-experts with political elites. To this end, we construct a bipartite network comprising of directed edges from PHEs/pseudo-experts to political elites with each edge indicating the political elite retweeted by a PHE or pseudo-expert. Figure 16(a) shows the distribution of ideology estimates for political elites active on Twitter during COVID-19. The ideology estimates are were obtained from 1 and were calculated using methods described in Barberá (2015). The median ideology score of political elites on Twitter is $-0.56$ which is consistent with the past reports identifying the liberal skew of Twitter Pew (2019). The median liberal and conservative elite have an ideology score of $-0.73$ and $1.28$ respectively. Interestingly, the median ideology score of the elites retweeted by PHEs and pseudo-experts are $-0.76$ and $1.54$ indicating that a considerable share of the elites retweeted by PHEs and pseudo-experts are more liberal and more conservative respectively, than the median liberal and conservative elite. In Figure 16(b) we visualize this network to find two highly clustered interactions. The color of the edge indicates the color of the source node i.e., PHEs (green) or pseudo-experts (orange). We find that PHEs mostly retweet liberal elites whereas, pseudo-experts mostly retweet conservative elites. This highlights the ideological clustering of the scientific elite in the United States.
@@
Figure 10(a) and (b) compares the impact of various emotions and moral foundations on engagement respectively. We find that, controlling for the number of followers, presence of anger and disgust in the original tweet generates more replies. This is also true for negative moral language: presence of harm, cheating, betrayal and subversion over its corresponding moral virtues is associated with more replies. These findings add important nuance to previous research showing that tweets expressing negative emotions Schöne et al. (2021) and moral outrage Brady et al. (2017) are more likely to be retweeted. Specifically, while negative emotions and moral language also receive more engagement in the form of replies, not all negative language leads to higher engagement: pessimism and sadness in the original tweets is associated with fewer replies. Importantly, positive language can even suppress engagement, as is the case for original PHE tweets expressing joy and love. Tables 6 and 7 tabulate the results of this regression analysis.
Figure 11: User reactions to (a) emotional and (b) moral appeals from PHEs. Demonstrates the odds ratio of users expressing particular emotions and moral principles in response to those conveyed in the original tweets by PHEs. An odds ratio greater than 1 signifies that the user is more inclined to express an emotion/moral foundation, equal to 1 indicates equivalent odds, and less than 1 suggests a decreased likelihood of expression. * indicates significance at $p<0.05$, ** - $p<0.01$, *** - $p<0.001$, **** - $p<0.0001$.
@@ Discussion
Firstly, we categorized tweets from PHEs and pseudo-experts to seven different issues—origins of the pandemic, stay-at-home lockdown mandates, masking mandates, healthcare infrastructure, reopening the education system, therapeutics and vaccinations. Our analysis of Twitter discourse from public health experts and pseudo-experts shows that they focused on different subsets of issues, similar to what was found in other work Nogara et al. (2022). While PHEs predominantly focused on promoting public health measures such as social distancing, masking, improving healthcare infrastructure, safer reopening of schools, pseudo-experts opposed lockdowns and mask mandates and promoted alternative views on therapeutics and virus origins.
@@
We note several areas of future work. Given that Twitter users are not a representative sample of the United States population, our findings may primarily reflect the perspectives of a specific demographic (i.e., younger, more liberal, better educated, more interested in politics) Wojcik and Hughes (2019). Future studies can instead focus on multiple platforms and incorporate multimodal data. While our study examines COVID-19-related discourse, there’s potential for investigations into scientific divisions in perspectives on polarized topics such as climate change and genetically modified foods. Moreover, exploring the growing debate on the factors contributing to the decline in adolescent mental health presents another avenue for inquiry.
Despite being state-of-the art models to identify emotions and moral language Alhuzali and Ananiadou (2021); Guo et al. (2023), the these models aren’t oracles. The advent of more powerful, albeit expensive, instruction-tuned language models like ChatGPT, future works can leverage them at scale to identify emotions and moral attitudes more accurately. However, given these tasks’ inherent ambiguity, this application might not be straightforward. We emphasize that the event related shifts in use of emotions and moral language allows us to make observational assertions rather than causal ones. Future work can attempt to run natural experiments to quantify the impact of events on different cohorts.Aside from this, the disruption in our access to Twitter’s Education Access API resulted in us only being able to collect replies for a subset of the PHEs tweets in our dataset. However, it is important to note that the subset of tweets we collected replies for were not intentionally sampled or biased in any way. Finally, our dataset, despite being extensive, is between January 2020 and January 2021, limiting our findings to this period and excluding potential shifts in perspectives that occurred later on.
@@ Conclusions
Author Contributions
AR, NS, SG, KL and LR designed the study and contributed to writing and editing the manuscript.
Acknowledgements
This work is supported in part by the National Science Foundation under grant CCF 2200256.
Conflicts of Interest
None declared.

### v2

@@ Results
The results presented in Supplementary Table LABEL:tab:perspec demonstrate the contrasting viewpoints between the two groups on various issues. Regarding the origins of the virus, PHEs generally lean towards the belief that it originated in a laboratory, albeit with some skepticism, while pseudo-experts heavily criticize China and its potential involvement in gain of function research. PHEs emphasize the importance of ongoing vigilance, adherence to stay-at-home orders, and widespread use of masks, whereas pseudo-experts question the effectiveness of lockdowns and mask mandates and criticize government intervention in these areas. On the topic of therapeutics, PHEs urge caution against self-prescribing drugs like Hydroxychloroquine, Azithromycin, and Ivermectin without evidence of their efficacy in treating COVID-19, whereas pseudo-experts advocate for the use of these medications.
@@
Emotions fluctuate over time and in response to events. Figure 6 illustrates the temporal dynamics of positive emotions expressed by PHEs and pseudo-experts. We leverage ChatGPT to summarize changes in emotions expressed and are discussed further in Supplementary Table LABEL:tab:emotions_summ. Optimism and joy among PHEs surge following the announcement of stay-at-home orders post March 15, 2020. This can be attributed to factors such as gratitude for guidance by then-New York Governor Andrew Cuomo and enhanced accessibility to COVID-19 testing. Similarly, we note a corresponding albeit smaller increase among pseudo-experts, particularly in response to President Trump’s management of the pandemic and France’s endorsement of HCQ as a viable COVID-19 treatment.
@@
We also examine the use of moral language by the two groups. Supplementary Figure 13 compares the distribution of the daily share of tweets expressing each moral foundation. Overall, PHEs use more positive moral language, emphasizing the dimensions of care, fairness, authority, loyalty and purity, while pseudo-experts tend to prefer the negative moral dimensions of harm, cheating, subversion and betrayal. The differences in use of moral language are more subdued compared to those for emotions. Supplementary Figure 14 illustrates the temporal dynamics of positive and negative moral language used by PHEs and pseudo-experts. We summarize the positive spikes using ChatGPT in Supplementary Table LABEL:tab:mft_summ.
@@
Figure 7 compares the log-odds along with the corresponding standard errors of estimation, to show which group used more emotional language to frame a specific issue. The plot highlights differences between the two groups and gives insights into emotionally charged issues. The biggest gap in emotions appears on the issue of lockdowns, where pseudo-experts are far more likely to express anger, disgust and sadness than PHEs. This position is consistent with the efforts to end the lockdowns (see Great Barrington declaration [94]). The second largest gap in emotions appears in the discussion of therapeutics, where PHEs are more likely to express anger and disgust, but pseudo-experts are less likely to use these emotions. Pseudo-experts also use more positive language with more joy and optimism when talking about therapeutics, in contrast to PHEs, consistent with the highly contentious debates about this issue. Other notable differences include pseudo-experts expressing more fear and less joy about vaccines, while PHEs express less fear and more optimism.
@@
Studies show that public response to the pandemic became polarized fairly quickly, with political partisanship shaping online activity and discussions about the pandemic already in the early stages of the pandemic [3, 21]. Online discussions also became emotionally polarized: when interacting with members of the opposite party, Twitter users expressed more anger and disgust, more toxicity, and less joy than in their interactions with same-party members [95]. Such interactions are characteristic of affective polarization [18], patterns of in-group love and out-group hate that have contributed to the growing partisan divide and the erosion of trust between the two parties in US. As a result, partisanship predicted the adoption of pandemic prevention measures more than other factors [96].
@@
Figure 15 lists the top ten accounts that are more likely to be mentioned by PHEs and pseudo-experts positively or negatively. We then assess the retweet interactions of public health and pseudo-experts with political elites. To this end, we construct a bipartite network comprising of directed edges from PHEs/pseudo-experts to political elites with each edge indicating the political elite retweeted by a PHE or pseudo-expert. Supplementary Figure 16(a) shows the distribution of ideology estimates for political elites active on Twitter during COVID-19. The ideology estimates are were obtained from 1 and were calculated using methods described in [97]. The median ideology score of political elites on Twitter is $-0.56$ which is consistent with the past reports identifying the liberal skew of Twitter [98]. The median liberal and conservative elite have an ideology score of $-0.73$ and $1.28$ respectively. Interestingly, the median ideology score of the elites retweeted by PHEs and pseudo-experts are $-0.76$ and $1.54$ indicating that a considerable share of the elites retweeted by PHEs and pseudo-experts are more liberal and more conservative respectively, than the median liberal and conservative elite. In Supplementary Figure 16(b) we visualize this network to find two highly clustered interactions. The color of the edge indicates the color of the source node i.e., PHEs (green) or pseudo-experts (orange). We find that PHEs mostly retweet liberal elites whereas, pseudo-experts mostly retweet conservative elites. This highlights the ideological clustering of the scientific elite in the United States.
@@
Figure 10(a) and (b) compares the impact of various emotions and moral foundations on engagement respectively. We find that, controlling for the number of followers, presence of anger and disgust in the original tweet generates more replies. This is also true for negative moral language: presence of harm, cheating, betrayal and subversion over its corresponding moral virtues is associated with more replies. These findings add important nuance to previous research showing that tweets expressing negative emotions [99] and moral outrage [11] are more likely to be retweeted. Specifically, while negative emotions and moral language also receive more engagement in the form of replies, not all negative language leads to higher engagement: pessimism and sadness in the original tweets is associated with fewer replies. Importantly, positive language can even suppress engagement, as is the case for original PHE tweets expressing joy and love. Supplementary Tables 6 and 7 tabulate the results of this regression analysis.
Figure 11: User engagement with PHEs and Pseudo Experts. User reactions to (a) emotional and (b) moral appeals from PHEs. Demonstrates the odds ratio of users expressing particular emotions and moral principles in response to those conveyed in the original tweets by PHEs. An odds ratio greater than 1 signifies that the user is more inclined to express an emotion/moral foundation, equal to 1 indicates equivalent odds, and less than 1 suggests a decreased likelihood of expression. * indicates significance at $p<0.05$, ** - $p<0.01$, *** - $p<0.001$, **** - $p<0.0001$.
@@ Discussion
Firstly, we categorized tweets from PHEs and pseudo-experts to seven different issues—origins of the pandemic, stay-at-home lockdown mandates, masking mandates, healthcare infrastructure, reopening the education system, therapeutics and vaccinations. Our analysis of Twitter discourse from public health experts and pseudo-experts shows that they focused on different subsets of issues, similar to what was found in other work [40]. While PHEs predominantly focused on promoting public health measures such as social distancing, masking, improving healthcare infrastructure, safer reopening of schools, pseudo-experts opposed lockdowns and mask mandates and promoted alternative views on therapeutics and virus origins.
@@
We note several areas of future work. Given that Twitter users are not a representative sample of the United States population, our findings may primarily reflect the perspectives of a specific demographic (i.e., younger, more liberal, better educated, more interested in politics) [100]. Future studies can instead focus on multiple platforms and incorporate multimodal data. While our study examines COVID-19-related discourse, there’s potential for investigations into scientific divisions in perspectives on polarized topics such as climate change and genetically modified foods. Moreover, exploring the growing debate on the factors contributing to the decline in adolescent mental health presents another avenue for inquiry.
Despite being state-of-the art models to identify emotions and moral language [22, 23], the these models aren’t oracles. The advent of more powerful, albeit expensive, instruction-tuned language models like ChatGPT, future works can leverage them at scale to identify emotions and moral attitudes more accurately. However, given these tasks’ inherent ambiguity, this application might not be straightforward. We emphasize that the event related shifts in use of emotions and moral language allows us to make observational assertions rather than causal ones. Future work can attempt to run natural experiments to quantify the impact of events on different cohorts.Aside from this, the disruption in our access to Twitter’s Education Access API resulted in us only being able to collect replies for a subset of the PHEs tweets in our dataset. However, it is important to note that the subset of tweets we collected replies for were not intentionally sampled or biased in any way. Finally, our dataset, despite being extensive, is between January 2020 and January 2021, limiting our findings to this period and excluding potential shifts in perspectives that occurred later on.